Net Present Value Analysis
Anderson Company must evaluate two capital expenditure proposals. Anderson's hurdle rate is 12%. Data for the two proposals follow.

Proposal X Proposal Y
Required investment $300,000 $300,000
Annual after-tax cash inflows 60,000
After-tax cash inflows at the end of years 3, 6, 9, and 12 180,000
Life of project 12 years 12 years

Using net present value analysis, which proposal is the more attractive?
Do not use negative signs with your answers. Round PV answers to the nearest whole number. Use rounded answers for subsequent calculation of net present value.

Proposal X Proposal Y
Net present value
Initial outflows Answer Answer
PV of future cash flows Answer Answer
Net present value Answer Answer

Which proposal is more attractive?
Answer #1
Proposal X Proposal Y
Net present value
Initial outflows $300,000 $300,000
PV of future cash flows 371662 330426
Net present value (71662) (30426)

Proposal Y is more attractive.

Working:

PV of future cash flows:

Proposal X = 60000*6.19437 = $371662

Proposal Y = 180000*0.71178+180000*0.50663+180000*0.36061+180000*0.25668 = $330426